A positive real number is 4 more than another. If the sum of the squares of the two numbers is 44, find the numbers.
Answer by jorel555(1290)

Let n be one of the numbers. Then the other number is n+4. So:
n²+(n+4)²=44
2n²+8n-28=0
n²+4n-14=0
Using the quadratic formula, we get two roots for n of 2.24264068712 and -6.24264068712. Throwing out the negative result, our real numbers are 2.24264068712 and 6.24264068712.